HUBBLE SPACE TELESCOPE DAILY REPORT #1719 PERIOD COVERED: 0000Z (UTC) 09/16/96 - 0000Z (UTC) 09/17/96 Daily Status Report as of 261/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 6906 (WF/PC-2 Cycle 6 Calibration: = Internal Monitor-Visual Flats) The WF/PC-2 was used to make internal calibration visual flat = images as part of the Cycle 6 internal monitoring of the instrument's = operations. This proposal defines the standard series of internal flat = fields, which will be executed twice per month. The observations were = completed as planned, and no anomalies were reported. 1.2 Completed FOS/RD 6796 (Spectral Mapping of Accretion Disks in = Cataclysmic Binaries) The Faint Object Spectrograph (Red Side) was used to make rapid = mode spectrographic mapping observations of the accretion disk around the = cataclysmic binary system V348-Puppis taken during the system's eclipse = phase. The observations were completed as planned, and no problems were = reported. 1.3 Completed Two Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science observati= ons. The intent is to search for extra-galactic sources with flux = excesses at 3000 A0 and galactic sources with excesses at 1600 A0. The = observations were completed as planned, and no problems were reported. 1.4 Completed WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS = Galaxy Sample-Part 1 High) The WF/PC-2 was used to make snapshot survey observations of the = ultraluminous IRAS galaxy QDOTULIR. T hese data will be searched for = multiple cores, which imply a merger has occurred, and for starburst = regions. The observations were executed as scheduled, and no anomalies = were noted. 1.5 Completed Two Sets of FOS/RD 6914 (FOS Cycle 6: Location of = Spectra and Wavelength Calibration - Red Side) The Faint Object Spectrograph (Red Side) was used to make both a = TALED and a wavelength calibration image. All observations were made = using internal light sources. This proposal will be executed a total of = 12 times during cycle 6 (i.e., once per month). The observations were = executed as scheduled, and no anomalies were noted. 1.6 Completed FOS/RD 6505 (Confirming the Quasar in Cygnus-A) The Faint Object Spectrograph (Red Side) was used to make = spectrographic observations of Cygnus-A in order to confirm the presence = of Mg II emission lines. Signs of this emission were seen in previous = HST observations, but there was a low signal/noise ratio and the = existence of an unflagged bad diode wade it necessary to acquire another = observation to determine if the feature was real. The observations were = executed as scheduled, and no anomalies were noted. 1.7 Completed Two Sets of WF/PC-2 7203 (Cosmology with the Medium Deep = Survey - Short Orbit Cases) ' The WF/PC-2 was used to take short (about 20-minute) exposures, to = be taken in parallel with the prime science observations, using the I = filter. These exposures are part of the medium-deep survey, which = emphasizes exposures that will be useful in obtaining cosmological = parameters. The observations were completed as planned, and no problems = were reported. 1.8 Completed FOS/RD 6589 (Associated Absorption Lines in QSOs) The Faint Object Spectrograph (Red Side) was used to make = spectrographic observations of the QSO PKS0237-23 in order to study its = associated absorption lines. The absorption lines (ZA) and the emission = lines (ZE) give similar redshifts, that is ZA~ZE. Recent studies have = suggested that the metallicity in its environment is higher than the = solar value. Near the beginning of this proposal, at 260/144618Z, a = NSSC-1 status buffer message was received, indicating that an observation = failed when no target was identified in the FOS aperture. HSTAR 5713 was = written. The effect on the remaining science is not known. However, the = remaining observations were completed as planned, and no further = anomalies were reported. 1.9 Completed WF/PC-2 6802 (Cosmology With the Medium Deep Survey) The WF/PC-2 was used to make observations taken in parallel with = the prime science observations as part of the Medium Deep Sky Survey. = The observations were executed as scheduled, and no problems were = reported. 1.10 Completed WF/PC-2 6348 (Blue Continuum Snapshots of 3CR Radio = Galaxies) The WF/PC-2 was used to make some snapshot images of the blue = continuum in radio galaxy 3C78.0. Combined with existing data, these = exposures will provide high spatial resolution information on the color = distribution within galaxies, redding due to dust, spacial indices of the = optical synchrotron jets, and colors of the active nuclei and tidal jets. = The observations were executed as scheduled, and no anomalies were = noted. 1.11 Completed FOS/BL 6498 (Accretion Streams in Magnetic Variables: = FOS Eclipse of 4 Rosa) The Faint Object Spectrograph (Blue Side) was used to make rapid = mode spectrographic observations of the accretion streams in the magnetic = variable star (i.e., cataclysmic variable) RXJ0515+01. This is a binary = system, and it contains a white dwarf star, an accretion disk, a mass = stream, and hot spot where the mass stream hits the disk and a companion = star. The observations were completed as planned and no anomalies were = reported. 1.12 Completed GHRS/2 6615 (Fundamental Parameters of White Dwarfs in = Close Binaries) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the close binary system EUEV-J0720 (a = white dwarf/M-dwarf binary system) in order to obtain fundamental = parameters of its white dwarf star. As the older star ages through the = asymptotic branch phase, the gaseous envelope which compasses both stars = dissipates and the white dwarf becomes visible. A wavelength calibration = image was also taken. The observations were executed as scheduled, and = no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled : 11 Successful : 11 2.3 Operations Notes: There were no reported operational problems. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH